# Robust Genomic Prediction and Heritability Estimation using Density Power Divergence

### v1

@@ Abstract
This manuscript delves into the intersection of genomics and phenotypic prediction, focusing on the statistical innovation required to navigate the complexities introduced by noisy covariates and confounders. The primary emphasis is on the development of advanced robust statistical models tailored for genomic prediction from single nucleotide polymorphism (SNP) data collected from genome-wide association studies (GWAS) in plant and animal breeding and multi-field trials. The manuscript explores the limitations of traditional marker-assisted recurrent selection, highlighting the significance of incorporating all estimated effects of marker loci into the statistical framework and aiming to reduce the high dimensionality of GWAS data while preserving critical information. This paper introduces a new robust statistical framework for genomic prediction, employing one-stage and two-stage linear mixed model analyses along with utilizing the popular robust minimum density power divergence estimator (MDPDE) to estimate genetic effects on phenotypic traits. The study illustrates the superior performance of the proposed MDPDE-based genomic prediction and associated heritability estimation procedures over existing competitors through extensive empirical experiments on artificial datasets and application to a real-life maize breeding dataset. The results showcase the robustness and accuracy of the proposed MDPDE-based approaches, especially in the presence of data contamination, emphasizing their potential applications in improving breeding programs and advancing genomic prediction of phenotyping traits.
Keywords: Genome-wide Association Study (GWAS), Genomic Prediction, Field Trials, Robust Estimation, Minimum Density Power Divergence Estimator, Linear Mixed Models.
@@ 1 Introduction
In the expansive landscape of statistical innovation, the intersection of genomics and phenotypic prediction constitutes a pivotal frontier. The ascent of high-throughput genotyping technologies, epitomized by the widespread adoption of single nucleotide polymorphism (SNP) data, has propelled a statistical revolution geared toward unraveling the genetic basis of complex traits. This manuscript unfolds as an answer to the escalating demand for advanced statistical models, meticulously tailored to exploit the richness of genomic data while navigating the intricacies introduced by covariates and confounders. In the realm of statistical genomics, genome-wide association studies (GWAS) serve as the vanguard and are a vital scientific method for identifying genetic variations associated with specific traits, diseases, or conditions in both animal and plant breeding (Uffelmann et al., 2021 [28]; Lourenço et al., 2020 [19]). This process is meticulously governed by an arsenal of statistical models and methods, each calibrated to distill meaningful insights from the vast genomic dataset. Unlike traditional marker-assisted recurrent selection, GWAS transcends the constraints of significance, incorporating all estimated effects of marker loci as well as genomic breeding values of individuals into its statistical framework to predict marker effects, irrespective of their individual association with the phenotype. The integration of genomics into phenotypic prediction assumes a pivotal role in deciphering the intricate statistical code governing trait variations. In the dynamic interplay of genomics and statistics, the significance of statistical models that transcend traditional limitations becomes paramount. Advanced statistical techniques enhance the power and precision of genetic association analyses, identifying subtle genetic effects often missed by traditional methods when dealing with numerous genetic markers. Many of these techniques aim to reduce high-dimensionality of GWAS data while preserving critical information, aiding in the identification of key genetic variants associated with traits (Giraud, 2021 [13]; Wainwright, 2019 [29]; Buhlmann et al., 2014 [5]). Some models, however, do take into account the high-dimensional nature of the marker matrix and attempt to fit the entire genome using methods such as penalized-likelihood or Bayesian shrinkage estimation. However, several issues have been identified when employing these models for effect estimation including computational intensity and modeling complexity. Shen et al. (2013) [26] proposed that utilizing heteroscedastic marker variances in ridge regression offers an alternative to Bayesian genome-wide prediction models. To illustrate the superior effectiveness of the heteroscedastic effect model, Hofheinz and Frisch (2014) [15] employed two methods within a one-stage model analysis where a single model incorporates all random effects (genetic effects, block effects, replicate effects, etc.) following classical literature. Alternatively, to avoid the computational complexity of this one-stage method, many authors proposed a second two-stage mixed model framework, where the random components are partitioned into two distinct stages; firstly genetic effects are considered as fixed effects while block and replicate effects are considered as random effects, and subsequently, in the second stage, genetic effects are regarded as random variables (Piepho et al., 2012 [21]; Estaghvirou et al., 2013 [6]). These types of data analyses have been commonly applied in plant and animal breeding techniques and multi-factor field trials in previous literature (Westhues et al., 2021 [30]; Tanaka, 2020 [27]; Resende, 2016 [24]; Bernal-Vasquez et al., 2016 [4]; Estaghvirou et al., 2015 [7]; Falke et al., 2014 [8]; Shen et al., 2013 [26]; Piepho, 2009 [22]; Möhring and Piepho, 2009 [20]; Barbosa et al., 2005[1]). This manuscript embarks on a statistical journey, navigating the challenges posed by noisy covariates and confounders to refine genomic prediction even under possible data contamination, aligning with the evolving demands of precision statistical analyses. We present a statistically robust framework for genomic prediction, building upon the foundation laid by GWAS. The overarching objectives revolve around the development and validation of a sophisticated robust statistical model adept at seamlessly integrating SNP data with a myriad of covariates. Through a statistically rigorous journey, we aim to showcase the model’s superiority in capturing the nuanced interplay between genetic markers, environmental factors, and covariate effects. The primary goal is to provide a powerful tool that not only enhances the precision and reliability of genomic predictions but also stands as a testament to the evolving landscape of statistical methodologies in genomics. Two distinct model based approaches are employed for this purpose, namely one-stage and two-stage linear mixed model analysis; in both the approaches we additionally propose to suitably utilize the popular robust minimum density power divergence estimator (MDPDE) to estimate genetic effects on various phenotypic traits. The MDPDE is a robust statistical procedure employed to address the complexities of the linear mixed model setup, emphasizing robustness and accuracy in obtaining reliable estimates of genetic effects (Lee and Jo, 2023 [18]; Saraceno et al., 2020 [25]; Goegebeur et al., 2014 [14]). Density power divergence (DPD) is pivotal to our analysis, quantifying the dissimilarity between probability density functions. It is a fundamental component of the robust estimation process, particularly suited for addressing contaminated data and model misspecification (Ghosh and Basu, 2013 [10]; Piepho et al., 2012b [21]). Notably, the MDPDE has gained widespread popularity due to its high robustness, computational simplicity, and its capacity to offer a direct interpretation akin to the maximum likelihood estimator (MLE). In our study, the robust statistical methods are deliberately designed to withstand influential factors, including the presence of outliers in various directions, deviations from normality, and issues stemming from model misspecification. To underscore the robustness of these techniques, we subject them to examination using Monte-Carlo simulations involving both uncontaminated and contaminated data. Furthermore, our investigation extends to practical application as we evaluate these methods by simulating data in two distinct setups and assessing their performance on a real dataset related to maize (Zea mays) genetics taken from Hofheinz and Frisch (2014) [15]. Additionally, in the present study, we propose a robust procedure for estimating heritability, as a by-product of genomic estimation, that leads to more accurate result than traditional heritability estimators especially when the data is contaminated with outliers. Heritability is unequivocally entangled to bioethics. To extent that twin heritability estimates played a central role in debates of science from the 70s through the 90s, the advent of SNP-heritability warrants a renewed assessment. It quantifies the proportion of phenotypic variation in a population that can be attributed to genetic factors (Zhu and Zhou, 2020 [32]), yielding an estimate of the extent to which genetic differences contribute to the observed variability in a particular trait within a population. We assess the effectiveness of the proposed robust estimator of heritability through Monte Carlo simulation studies and by applying them to an actual maize dataset mentioned above.
@@ 2.1 Mathematical Model for Genomic Prediction
Suppose that our data consist of $N$ individual phenotype observations in $r$ replicate and $p$ biallelic single-nucleotide polymorphism (SNP) markers where genotype effects are distributed in $B$ blocks and $L$ confounder or controllers. Let $\mathbf{y}_{k}$ denotes the vector of $n_{k}$ phenotype data for the $k$-th replicate for each $k=1,\ldots,r$, $\mathbf{X}_{g}$ denotes the matrix of $p$ SNP markers along the genome for $N$ individuals ($\mathbf{X}_{g}$ has $p$ columns for the SNP observations, usually coded as 0, 1, --1 for homo zygote aa, the hetero-zygote Aa and the other homo-zygote AA, respectively) $\mathbf{X}_{b}$ denotes the design matrix for random block effect and $\mathbf{Z}$ denotes the $N\times L$ matrix of confounders. We denote the $k$-th partition of rows of $\boldsymbol{Z}$, $\mathbf{X}_{g}$ and $\mathbf{X}_{b}$ associated with each $\mathbf{y}_{k}$ by $\boldsymbol{Z}_{k}$, $\boldsymbol{X}_{gk}$ and $\boldsymbol{X}_{bk}$, respectively, for all $k=1,\ldots,r$. Following preceding works, e.g., Tanaka, 2020 [27]; Lourenço et al., 2020 [19]; Estaghvirou et al., 2014 [6]; Piepho et al., 2012 [23], let us assume the a linear mixed model (LMM) for SNP data in the $k$-th replicate as given by
$\mathbf{y}_{k}=\mathbf{Z}_{k}\boldsymbol{\gamma}+\mathbf{X}_{gk}\mathbf{u}_{g}+\mathbf{X}_{bk}\mathbf{u}_{b}+\boldsymbol{\epsilon}_{k},~{}~{}~{}~{}k=1,\ldots,r,$ (1)
where the $p$-vector of random SNP effect $\mathbf{u}_{g}\sim N(\mathbf{0},\boldsymbol{\Sigma}_{g})$, the random block effects $\mathbf{u}_{b}\sim N(\mathbf{0},\sigma_{b}^{2},\mathbb{I}_{B})$, $\boldsymbol{\gamma}$ denotes the $l$-vector of the confounder effects and the random error $\boldsymbol{\epsilon}_{k}\sim N(\boldsymbol{0},\sigma_{e}^{2}\mathbb{I}_{n_{k}})$ for $k=1,\ldots,r$. Note that $N=\sum_{k=1}^{r}n_{k}$ which is the total number of observation and individual replication is assumed to be independent of each other. Put $\boldsymbol{y}=(\boldsymbol{y}^{{}^{\prime}}_{1},\boldsymbol{y}^{{}^{\prime}}_{2},....,\boldsymbol{y}^{{}^{\prime}}_{r})^{{}^{\prime}}$ to denote the vector of all phenotype data. Then, we can rewrite the LMM in (1) as
$\mathbf{y}=\mathbf{Z}\boldsymbol{\gamma}+\begin{bmatrix}\mathbf{X}_{g}&\mathbf{X}_{b}\end{bmatrix}\begin{bmatrix}\mathbf{u}_{g}\\ \mathbf{u}_{b}\\ \end{bmatrix}+\boldsymbol{\epsilon}=\mathbf{Z}\mathbf{\gamma}+\mathbf{X}\mathbf{u}+\boldsymbol{\epsilon},$ (2)
@@ 2.2 Estimating Marker Effects and Breeding Values
$\begin{bmatrix}{\textbf{Z}}^{T}\mathbf{Z}&\textbf{Z}^{T}\mathbf{X}\\ \textbf{X}^{T}\mathbf{Z}&\textbf{X}^{T}\mathbf{X}+\mathbf{\Lambda}\end{bmatrix}\begin{bmatrix}\boldsymbol{\gamma}\\ \textbf{u}\end{bmatrix}=\begin{bmatrix}\textbf{Z}^{T}\mathbf{y}\\ \textbf{X}^{T}\mathbf{y}\end{bmatrix},$ (3)
where $\boldsymbol{\Lambda}=Diag(\lambda_{1},\lambda_{2},.....,\lambda_{p})$, with $\lambda_{j}$ being the shrinkage parameter of $j$-th SNP for $j=1,\ldots,p$, which is indeed taken as a function of $\sigma_{e}^{2}$ and $\boldsymbol{\Sigma}_{g}$. For estimating the marker-effects, the traditional approach is to minimize the likelihood based loss function associated with the model (2) which can be simplified to have the form
$L(\mathbf{u},\boldsymbol{\gamma})=(\mathbf{y}-\mathbf{Xu}-\mathbf{Z}\gamma)^{T}(\mathbf{y-Xu-Z}\gamma)+\mathbf{u}^{T}\Lambda\mathbf{u}.$
This loss function can be minimized analytically to get an estimator of the random-effect $u$ (including both SNP-marker effects and block effects) as given by
$\widehat{\mathbf{u}}=(\mathbf{X}^{T}\mathbf{M_{Z}X}+\boldsymbol{\Lambda})^{-1}\mathbf{X}^{T}\mathbf{M_{Z}y},$ (4)
where $\mathbf{M_{Z}=I-Z(Z}^{T})^{-1}\mathbf{Z}^{T}$ is the projection matrix removing the effect of confounding or controller variables, $\boldsymbol{\Lambda}$ is a diagonal matrix of shrinkage parameters $\lambda_{j}$, $j=1,\ldots,p$. These shrinkage parameters $\lambda_{j}$, being a function of the variance parameters, can be estimated from different types of variance estimators and subsequently we can get the estimated $\widehat{\mathbf{u}}$ from (4) using estimated values of $\lambda_{j}$s in $\boldsymbol{\Lambda}$. Then, for genomic prediction, the estimate of breeding value will be given by [26]
@@
We would like to emphasis that there exist several approaches for obtaining the estimates of random effects and error variances leading to different estimates of the shrinkage parameters in the literature of genomic prediction. For ordinary ridge regression, we assume the simplest form $\boldsymbol{\Sigma}_{g}=\sigma_{g}^{2}\mathbb{I}_{p}$ and then term $\boldsymbol{\Lambda}$ can be replaced by $\lambda\mathbb{I}$ with the constant shrinkage parameter being computed as $\widehat{\lambda}=\frac{\widehat{\sigma}_{e}^{2}}{\widehat{\sigma}_{g}^{2}}$. However, while using the likelihood based estimation approaches, it has been shown that the heteroscedastic values of $\lambda_{j}$ works significantly better where we assume $\boldsymbol{\Sigma}_{g}=Diag\left(\sigma_{g1}^{2},\ldots,\sigma_{gp}^{2}\right)$ and estimate different shrinkage parameters, corresponding to different SNPs, as $\widehat{\lambda}_{j}=\frac{\widehat{\sigma}_{e}^{2}}{\widehat{\sigma}_{gj}^{2}}$ for every $j=1,\ldots,p$. The most common one is to use the restricted maximum likelihood estimator under the assumed LMM which is often referred to as the ridge regression (RR) estimator. Hofheinz and Frisch (2014) [15] proposed two methods, to be referred to as RMLV and RMLA, that are shown to have significantly improved performances over the classical RR estimator for genomic prediction; they both estimate the random effect $\widehat{\mathbf{u}}$ using (4) with $\lambda_{j}$ replaced by its estimators $\widehat{\lambda}_{j}$ for all $j$ under the model assumption $\boldsymbol{\Sigma}_{g}=Diag\left(\sigma_{g1}^{2},\ldots,\sigma_{gp}^{2}\right)$, but these $\widehat{\lambda}_{j}$ are obtained in two ways as described below:
@@
    $\widehat{\sigma}_{e}^{2}=(\textbf{y}^{{}^{\prime}}\textbf{y}-\widehat{\boldsymbol{\gamma}^{{}^{\prime}}}\textbf{Z}^{{}^{\prime}}\mathbf{y}-\widehat{\textbf{u}^{{}^{\prime}}}X^{{}^{\prime}}\textbf{y})/(N-1),~{}~{}~{}\mbox{ and }~{}~{}~{}\widehat{\sigma}_{gj}^{2}=(\widehat{\textbf{u}_{j}}^{{}^{\prime}}\widehat{\mathbf{u}_{j}}-\sigma_{e}^{2}\cdot tr(\mathbf{C}_{jj}))/j,$

### v2

@@ Abstract
This manuscript delves into the intersection of genomics and phenotypic prediction, focusing on the statistical innovation required to navigate the complexities introduced by noisy covariates and confounders. The primary emphasis is on the development of advanced robust statistical models tailored for genomic prediction from single nucleotide polymorphism data in plant and animal breeding and multi-field trials. The manuscript highlights the significance of incorporating all estimated effects of marker loci into the statistical framework and aiming to reduce the high dimensionality of data while preserving critical information. This paper introduces a new robust statistical framework for genomic prediction, employing one-stage and two-stage linear mixed model analyses along with utilizing the popular robust minimum density power divergence estimator (MDPDE) to estimate genetic effects on phenotypic traits. The study illustrates the superior performance of the proposed MDPDE-based genomic prediction and associated heritability estimation procedures over existing competitors through extensive empirical experiments on artificial datasets and application to a real-life maize breeding dataset. The results showcase the robustness and accuracy of the proposed MDPDE-based approaches, especially in the presence of data contamination, emphasizing their potential applications in improving breeding programs and advancing genomic prediction of phenotyping traits.
Keywords: Genomic Prediction, Field Trials, Robust Estimation, Minimum Density Power Divergence Estimator, Linear Mixed Models.
@@ 1 Introduction
In the expansive landscape of statistical innovation, the intersection of genomics and phenotypic prediction constitutes a pivotal frontier. The rise of high-throughput genotyping technologies, such as the extensive utilization of single nucleotide polymorphism (SNP) data, is driving a significant shift in statistical methodologies towards revealing the genetic foundations of intricate traits. The present manuscript contributes to the growing need for sophisticated statistical models tailored to employ the richness of genomic data while navigating the complexities introduced by covariates, confounders, and different noises (contamination) in the data, specifically with the objective of genomic prediction (of phenotypic traits) using data from randomly sampled individuals from a population.
In the realm of statistical genomics, genome-wide association studies serve as an important scientific method for identifying genetic variations associated with specific traits, diseases, or conditions in both plant and animal breeding (Uffelmann et al., 2021 [32]; Lourenço et al., 2020 [21]). This process relies on a plethora of statistical models and methods, each crafted to refine meaningful insights from vast genomic datasets. Unlike traditional marker-assisted recurrent selection, genomic prediction integrates all estimated effects of marker loci to predict genomic breeding values of individuals, regardless of their individual association with the phenotype. Advanced statistical techniques enhance the accuracy of such genomic prediction, uncovering the acute genetic effects often overlooked by conventional methods in handling extensive genetic markers. Many of these techniques aim to reduce high-dimensionality while preserving critical information (Giraud, 2021 [17]; Wainwright, 2019 [33]; Buhlmann et al., 2014 [5]). Some models, however, do take into account the high-dimensional nature of the marker matrix and attempt to fit the entire genome using methods such as penalized-likelihood or Bayesian shrinkage estimation. However, several issues have been identified when employing these models for effect estimation including computational intensity and modeling complexity. Shen et al. (2013) [30] suggested using heteroscedastic marker variances in ridge regression as an alternative to Bayesian genome-wide prediction models. Hofheinz and Frisch (2014) [18] demonstrated the superior effectiveness of the heteroscedastic effect model by employing two methods within a one-stage linear mixed model analysis, encompassing all random effects (genetic effects, block effects, replicate effects, etc.) following classical literature. Alternatively, to avoid the computational complexity of this one-stage method, many authors proposed a second two-stage mixed model framework, where the random components are partitioned into two distinct stages; firstly genetic effects are considered as fixed effects while block and replicate effects are considered as random effects, and subsequently, in the second stage, genetic effects are regarded as random effects (Piepho et al., 2012 [24]; Estaghvirou et al., 2013 [8]). These types of mixed models have been commonly applied in plant breeding techniques and multi-environment trials in previous literature (Westhues et al., 2021 [34]; Tanaka, 2020 [31]; Resende, 2016 [27]; Bernal-Vasquez et al., 2016 [4]; Estaghvirou et al., 2015 [9]; Falke et al., 2014 [10]; Shen et al., 2013 [30]; Piepho, 2009 [25]; Möhring and Piepho, 2009 [22]; Barbosa et al., 2005[1]).
In the present study, we develop a new robust statistical procedure for genomic prediction under possible data contamination, based on complex noisy data, effectively capturing complex interactions among genetic markers, environmental factors, and covariates. For these, we employ two model-based approaches, namely the one-stage and two-stage linear mixed model analyses from the literature of genomic prediction, but propose to use the popular robust minimum density power divergence (DPD) estimation instead of classical maximum likelihood based estimation procedure. The minimum DPD estimator (MDPDE), after originally introduced by Basu et al. (1998) [2], has subsequently gained widespread popularity due to its strong robustness properties with just a little loss in (asymptotic) efficiency, computational simplicity, and its capacity to offer a direct interpretation akin to the maximum likelihood estimator (MLE); see, e.g., Basu et al. (2011) [3], Ghosh and Basu (2013) [14] and the references therein. Saraceno et al. (2020) [28] have recently studied the MDPDE for linear mixed effect models, which we employ here for developing robust genomic prediction under one and two stage mixed model approaches. As a result, our proposed methods withstand influential factors, including the presence of outliers in various directions and similar issues stemming from model misspecification, providing improved performances over the latest existing procedures of genomic prediction, as illustrated through extensive Monte-Carlo simulations with both uncontaminated and contaminated data. Furthermore, we evaluate applicability and performances of these proposed methods with a real dataset of a maize (Zea mays) breeding study, taken from Hofheinz and Frisch (2014) [18].
Additionally, in the present study, we propose a robust procedure for estimating heritability, as a by-product of genomic prediction, that leads to more accurate result than traditional heritability estimators especially in the presence of outliers (or other noises/contamination) in the data. Heritability is inherently entangled to bioethics. The advent of SNP-heritability warrants a renewed assessment, quantifying the genetic contribution to phenotypic variation within a population (Zhu and Zhou, 2020 [35]). We assess the effectiveness of the proposed robust heritability estimator also through Monte Carlo simulation studies and by applying them to an actual maize dataset mentioned above.
@@ 2.1 Mathematical Model for Genomic Prediction
Suppose that our data consist of $N$ individual phenotype observations in $r$ replicate and $p$ biallelic single-nucleotide polymorphism (SNP) markers where genotype effects are distributed in $B$ blocks and $L$ confounder or controllers in a randomized complete block design. The block often represents different crop types or soil/environmental conditions in plant breeding and multi-environment trials, and confounding/controlling variables are any additional covariates (e.g., replication indicator, surrounding temperature/humidity, soil properties, etc., when they are not used to construct the blocks) that we wish to include in our analyses as random effects from subject-matter knowledge on the actual experimental setting. Mathematically, let $\boldsymbol{y}_{k}$ denotes the vector of $n_{k}$ phenotype data for the $k$-th replicate for each $k=1,\ldots,r$, $\mathbf{X}_{g}$ denotes the matrix of $p$ SNP markers along the genome for $N$ individuals ($\mathbf{X}_{g}$ has $p$ columns for the SNP observations, usually coded as 0, 1, --1 for homo zygote aa, the hetero-zygote Aa and the other homo-zygote AA, respectively) $\mathbf{X}_{b}$ denotes the design matrix for random block effect and $\mathbf{Z}$ denotes the $N\times L$ matrix of confounders. We denote the $k$-th partition of rows of $\boldsymbol{Z}$, $\mathbf{X}_{g}$ and $\mathbf{X}_{b}$ associated with each $\boldsymbol{y}_{k}$ by $\boldsymbol{Z}_{k}$, $\boldsymbol{X}_{gk}$ and $\boldsymbol{X}_{bk}$, respectively, for all $k=1,\ldots,r$. Following preceding works, e.g., Tanaka (2020) [31], Lourenço et al. (2020) [21], Estaghvirou et al. (2014) [8], and Piepho et al. (2012) [26], let us assume the a linear mixed model (LMM) for SNP data in the $k$-th replicate as given by
$\boldsymbol{y}_{k}=\mathbf{Z}_{k}\boldsymbol{\gamma}+\mathbf{X}_{gk}\mathbf{u}_{g}+\mathbf{X}_{bk}\mathbf{u}_{b}+\boldsymbol{\epsilon}_{k},~{}~{}~{}~{}k=1,\ldots,r,$ (1)
where the $p$-vector of random SNP effect $\mathbf{u}_{g}\sim N(\mathbf{0},\boldsymbol{\Sigma}_{g})$, the random block effects $\mathbf{u}_{b}\sim N(\mathbf{0},\sigma_{b}^{2},\mathbb{I}_{B})$, $\boldsymbol{\gamma}$ denotes the $l$-vector of the fixed effects and the random error $\boldsymbol{\epsilon}_{k}\sim N(\boldsymbol{0},\sigma_{e}^{2}\mathbb{I}_{n_{k}})$ for $k=1,\ldots,r$. Note that $N=\sum_{k=1}^{r}n_{k}$ which is the total number of observation and individual replication is assumed to be independent of each other; the data are partitioned by replicates to utilize this independence across replications in the mathematical formulation of our robust estimation procedures.
Putting $\boldsymbol{y}=(\boldsymbol{y}^{{}^{\prime}}_{1},\boldsymbol{y}^{{}^{\prime}}_{2},....,\boldsymbol{y}^{{}^{\prime}}_{r})^{{}^{\prime}}$ to denote the vector of all phenotype data, we can rewrite the LMM in (1) as
$\boldsymbol{y}=\mathbf{Z}\boldsymbol{\gamma}+\begin{bmatrix}\mathbf{X}_{g}&\mathbf{X}_{b}\end{bmatrix}\begin{bmatrix}\mathbf{u}_{g}\\ \mathbf{u}_{b}\\ \end{bmatrix}+\boldsymbol{\epsilon}=\mathbf{Z}\mathbf{\gamma}+\mathbf{X}\mathbf{u}+\boldsymbol{\epsilon},$ (2)
@@ 2.2 Estimating Marker Effects and Breeding Values
$\begin{bmatrix}{\textbf{Z}}^{T}\mathbf{Z}&\textbf{Z}^{T}\mathbf{X}\\ \textbf{X}^{T}\mathbf{Z}&\textbf{X}^{T}\mathbf{X}+\mathbf{\Lambda}\end{bmatrix}\begin{bmatrix}\boldsymbol{\gamma}\\ \textbf{u}\end{bmatrix}=\begin{bmatrix}\textbf{Z}^{T}\boldsymbol{y}\\ \textbf{X}^{T}\boldsymbol{y}\end{bmatrix},$ (3)
where $\boldsymbol{\Lambda}=Diag(\lambda_{1},\lambda_{2},.....,\lambda_{p})$, with $\lambda_{j}$ being the shrinkage parameter of $j$-th SNP for $j=1,\ldots,p$, which is indeed taken as a function of $\sigma_{e}^{2}$ and $\boldsymbol{\Sigma}_{g}$.
For estimating the marker-effects, the traditional approach is to minimize the likelihood based loss function associated with the model (2) which can be simplified to have the form
$L(\mathbf{u},\boldsymbol{\gamma})=(\boldsymbol{y}-\mathbf{Xu}-\mathbf{Z}\gamma)^{T}(\mathbf{y-Xu-Z}\gamma)+\mathbf{u}^{T}\Lambda\mathbf{u}.$
This loss function can be minimized analytically to get an estimator of the random-effect $\mathbf{u}$ (including both SNP-marker effects $\mathbf{u}_{b}$ and block effects $\mathbf{u}_{b}$) as given by
$\widehat{\mathbf{u}}=\begin{bmatrix}\widehat{\mathbf{u}}_{g}\\ \widehat{\mathbf{u}}_{b}\end{bmatrix}=(\mathbf{X}^{T}\mathbf{M_{Z}X}+\boldsymbol{\Lambda})^{-1}\mathbf{X}^{T}\mathbf{M_{Z}y},$ (4)
where $\mathbf{M_{Z}}=\mathbb{I}_{N}-\mathbf{Z}(\mathbf{Z}^{T}\mathbf{Z})^{-1}\mathbf{Z}^{T}$ is the projection matrix removing the effect of confounding or controller variables, $\boldsymbol{\Lambda}$ is a diagonal matrix of shrinkage parameters $\lambda_{j}$, $j=1,\ldots,p$. These shrinkage parameters $\lambda_{j}$, being a function of the variance parameters, can be estimated from different types of variance estimators and subsequently we can get the estimated $\widehat{\mathbf{u}}$ from (4) using estimated values of $\lambda_{j}$s in $\boldsymbol{\Lambda}$. Then, for genomic prediction, the estimate of breeding value will be given by [30]
@@
We would like to emphasize that there exist several approaches for obtaining the estimates of random effects and error variances leading to different estimates of the shrinkage parameters in the literature of genomic prediction. For ordinary ridge regression, we assume the simplest form $\boldsymbol{\Sigma}_{g}=\sigma_{g}^{2}\mathbb{I}_{p}$ and then term $\boldsymbol{\Lambda}$ can be replaced by $\lambda\mathbb{I}$ with the constant shrinkage parameter being computed as $\widehat{\lambda}=\frac{\widehat{\sigma}_{e}^{2}}{\widehat{\sigma}_{g}^{2}}$. However, while using the likelihood based estimation approaches, it has been shown that the heteroscedastic values of $\lambda_{j}$ works significantly better where we assume $\boldsymbol{\Sigma}_{g}=Diag\left(\sigma_{g1}^{2},\ldots,\sigma_{gp}^{2}\right)$ and estimate different shrinkage parameters, corresponding to different SNPs, as $\widehat{\lambda}_{j}=\frac{\widehat{\sigma}_{e}^{2}}{\widehat{\sigma}_{gj}^{2}}$ for every $j=1,\ldots,p$. The most common one is to use the restricted maximum likelihood estimator under the assumed LMM which is often referred to as the ridge regression (RR) estimator. Hofheinz and Frisch (2014) [18] proposed two methods, to be referred to as RMLV and RMLA, that are shown to have significantly improved performances over the classical RR estimator for genomic prediction; they both estimate the random effect $\widehat{\mathbf{u}}$ using (4) with $\lambda_{j}$ replaced by its estimators $\widehat{\lambda}_{j}$ for all $j$ under the model assumption $\boldsymbol{\Sigma}_{g}=Diag\left(\sigma_{g1}^{2},\ldots,\sigma_{gp}^{2}\right)$, but these $\widehat{\lambda}_{j}$ are obtained in two ways as described below:
@@
    $\widehat{\sigma}_{e}^{2}=(\textbf{y}^{{}^{\prime}}\textbf{y}-\widehat{\boldsymbol{\gamma}^{{}^{\prime}}}\textbf{Z}^{{}^{\prime}}\boldsymbol{y}-\widehat{\textbf{u}^{{}^{\prime}}}X^{{}^{\prime}}\textbf{y})/(N-1),~{}~{}~{}\mbox{ and }~{}~{}~{}\widehat{\sigma}_{gj}^{2}=(\widehat{\textbf{u}_{j}}^{{}^{\prime}}\widehat{\mathbf{u}_{j}}-\sigma_{e}^{2}\cdot tr(\mathbf{C}_{jj}))/j,$